Light with a wavelength of 600.0 nm is directed at a metallic surface with a work function of 0.96 eV. Find the max kinetic energy and max speed of the photoelectronsWhat is the cutoff potential necessary to stop the photoelectrons? 


1
Expert's answer
2021-04-19T17:07:20-0400

According to the Einstein's formula (see https://en.wikipedia.org/wiki/Photoelectric_effect#Theoretical_explanation) the max kinetic energy of the released electrons will be:



"K = h\nu - W"

where "W = 0.98eV" is the work function of the metal, "h = 4.1\times 10^{-15}eV\cdot s" is the Planck constant, and "\nu" is the frequency of the light. The frequency is connected wiht wavelength "\lambda = 600nm = 6\times 10^{-7}m" as follows:



"\nu = \dfrac{c}{\lambda}"

where "c = 3\times 10^{8}m/s" is the speed of light. Subsittuting this into the expression for "K", obtain:



"K = \dfrac{hc}{\lambda} - W"


Substituting the numbers, find:


"K = \dfrac{4.1\times 10^{-15}\cdot 3\times 10^8}{6\times 10^{-7}} - 0.98 = 1.07eV \approx 1.6\times 10^{-19}J"


On the other hand, the kinetic energy is given by the following formula (assuming non-relativistic case):


"K = \dfrac{mv^2}{2}"

where "m = 9.11\times 10^{-31}kg" is the mass of electron, and "v" is its maximum speed. Expressing "v", obtain:


"v = \sqrt{\dfrac{2K}{m}} = \sqrt{\dfrac{2\cdot 1.6\times 10^{-19}J}{9.11\times 10^{-31}kg}}\approx 5.93\times 10^5m/s"


In order to completely reduce this kinetic energy one should apply the following potential difference (according to the energy-work theorem):


"V = K"

if "K" is measured in eV. Then "V" will be in volts. Thus, obtain:


"V = 1.07V"



Answer. Max kinetic energy: "1.07eV", max speed: "5.93\times 10^5m/s", stop potential: "1.07V".